# Fractional Chern Insulator and Quantum Anomalous Hall Crystal in Twisted MoTe2

### v2

@@ III.1 Global phase diagram
Since the presence of FM order has been experimentally identified as a prerequisite for the emergence of FCI and other exotic topological states in MoTe2 [7, 10, 9, 8, 12, 11, 13], we first examine the emergence of FM order, characterized by full polarization of spin-$S_{z}$ (denoted as FMz). To evaluate the ground-state magnetism across various filling factors, we compare the energies of different $S_{z}$ sectors in our iDMRG simulations (see Supplemental Sec. I [56]). Our results reveal a finite energy gap above the FMz ground state for $\nu\gtrsim 0.2$, as indicated by darker blue shading in Fig. 1d for $N_{y}=4$ and $5$. Additionally, for $N_{y}=6$, we find that FMz order remains robust at filling factors $\nu=1$, $2/3$, $1/2$, and $1/3$, consistent with recent experimental observations [7, 8].
@@
To investigate the presence of quantized Hall conductivity in the FMz phases, we employ a charge pumping method [69, 70, 71] within our iDMRG simulations. As depicted in the inset of Fig. 2g, for filling factors $\nu=p/q$, we adiabatically introduce a flux of $\Phi_{y}=2q\pi$ along the cylinder axis by modulating the hopping terms with a phase factor $e^{i\Phi_{y}}c^{\dagger}_{i\sigma}c_{j\sigma}+h.c.$ as the electron traverses the periodic boundary. For a charge pumping of $\Delta Q$, the corresponding Hall conductivity is given by $\sigma_{xy}=\frac{2\pi|\Delta Q|}{\Phi_{y}}$, whose unit is set as $e^{2}/h$ in this study.
@@
Our simulations, primarily conducted for fillings $\nu=p/q$ and $q=1,2,3$, reveal the existence of quantized Hall conductivity $\sigma_{xy}$, as indicated in Fig. 1e. We also consider $q=5,7,9$ cases on $N_{y}=4,5$ cylinders, with the results summarized in Supplementary Fig. S8 [56]. Notably, we successfully observe experimentally relevant $\sigma_{xy}=\nu=1$ for CI states [56] and $\sigma_{xy}=\nu=2/3$, $4/7$, $5/9$, $4/9$, $3/5$ for FCIs. Intriguingly, we predict a robust FCI state with $\sigma_{xy}=\nu=1/3$ and advocate for its experimental observation in future studies. Moreover, we also identify instances where an integer Hall conductivity of $\sigma_{xy}=1$ can be observed at fractional fillings $\nu=1/2$, $2/3$ (for $N_{y}=6$ cylinder), as well as at $\nu=3/5$ and $4/5$ (for $N_{y}=5$). These integer Hall states coexist with COs that are compatible with the cylinder geometry, leading to the formation of QAHC states.
@@ III.2 GWC and FCI for $\nu=1/3$
To start, we consider the filling fraction $\nu=1/3$, and present the results on $N_{y}=5$ and $6$ cylinders. Our results reveal that the ground states exhibit FMz order for a broad range of $\epsilon_{r}$ values, specifically $10\leq\epsilon_{r}\leq 23$. For the regime with a relatively small $\epsilon_{r}$, a GWC phase appear in the phase diagram Fig. 2a. The real-space charge density $n(\bm{r})$ is centered at the $A$ sublattice and forms a triangular lattice with primitive vectors $\bm{a}^{t}_{1}=(\sqrt{3},0)$ and $\bm{a}^{t}_{2}=(\sqrt{3}/2,3/2)$, thus breaking the translational symmetry of the original honeycomb lattice (see Fig. 2d). Correspondingly, in Fig. 2e we find clear peaks in $\rho(\bm{q})$ at the vertices ($\bm{K}$ points), namely, the characteristic reciprocal vectors $\bm{b}^{t}_{1}=(0,4\pi/3)$ and $\bm{b}^{t}_{2}=(-2\sqrt{3}\pi/3,2\pi/3)$ of the triangular lattice. The results confirm the presence of a triangular GWC for relatively small $\epsilon_{r}$.
Figure 2: GWC and FCI phases for $\nu=1/3$ filling. a Phase diagram containing the GWC and FCI states tuned by the relative dielectric constant $\epsilon_{r}$. The two phases are characterized by $\rho({\bm{q}})$ and the Hall conductivity $\sigma_{xy}$, obtained on $N_{y}=5$ and $6$ cylinders. b Charge-charge correlation function $g^{nn}(\bm{r})$ in the FCI phase (with $\epsilon_{r}=18$), displaying the charge distribution relative to the reference site $n_{0}$ marked by a black star. c Charge structure factor $S^{A}_{n}(\bm{q})$ in momentum space of the FCI state. Charge distribution d $n(\bm{r})$ in real-space and e $\rho(\bm{q})$ in momentum space for the GWC state (with $\epsilon_{r}=14$), showcasing the triangular lattice pattern of electron density modulation. f demonstrates $n(\bm{r})$ for the FCI state ($\epsilon_{r}=18$), where a uniform charge distribution is observed. g Charge pumping results $|\Delta Q|$ for various $\epsilon_{r}$ for $N_{y}=6$, measured after inserting a flux $\Phi_{y}=6\pi$. The inset illustrates the charge pumping simulation process. For this and subsequent plots, the dashed hexagon (c.f., panels c and e) outlines the first Brillouin zones of the honeycomb lattice.
@@
However, for a larger relative dielectric constant, e.g., $\epsilon_{r}=18$, the charge density $n(\bm{r})$ becomes uniformly distributed across the honeycomb lattice (Fig. 2f). As a result, $\rho(\bm{q})$ becomes negligible across the entire momentum space (Fig. 2a). As seen in Fig. 2a, the maximum of $\rho(\bm{q})$ undergoes a sharp drop to very small values at $\epsilon^{c}_{r}=15.5$. For even larger $\epsilon_{r}>23$, the ground state no longer has an FMz order.
@@
To characterize the transitions between GWC and FCI, we compute the charge-charge correlations $g^{nn}(\bm{r})=\frac{\langle n_{0}n_{i}\rangle}{\langle n_{0}\rangle\langle n_{i}\rangle}$ between the charge density $n_{0}$ at a fixed site and other charge densities $n_{i}$ located at site $i$. This metric provides insights into the attractive ($g^{nn}>1$) and repulsive ($g^{nn}<1$) correlations between electrons. As depicted in Fig. 2b for $\epsilon_{r}=18$, we find $g^{nn}<1$ for the first three nearest neighbors of the reference site, indicating a repulsive interaction among electrons within this close proximity. However, $g^{nn}$ exceeds 1.0 for the fourth and fifth nearest neighbors, suggesting an attractive correlation in this region. This oscillating pattern of repulsion and attraction can extend even further away from the reference site, which constitutes a characteristic of neutral collective excitation, roton, in a FCI liquid [72].
@@
which quantifies the spatial distribution of charge correlations in momentum space. In Fig. 2c, we plot $S^{A}_{n}(\bm{q})$ for $\epsilon_{r}=18$ in the uniform state with $\rho^{m}(\bm{q})\simeq 0$. The $S^{A}_{q}(\bm{q})$ peaks at the $\bm{K}$ points, which reflect the oscillating pattern observed in $g^{nn}(\bm{r})$. The positions of the peaks align with those of the GWC peaks depicted in Fig. 2e, indicating that the condensation of roton modes may be responsible for the transition from a uniform FCI to a triangular-lattice GWC.
@@
Moreover, we simulate the Hall conductivity through charge pumping, as illustrated in Fig. 2g for several values of $\epsilon_{r}$ and $N_{y}=6$, using iDMRG calculations. Our results demonstrate that inserting a flux of $\Phi_{y}=6\pi$ induces a charge pumping of $|\Delta Q|=1$ (thus $\sigma_{xy}=1/3$) for FCIs, while $|\Delta Q|=0$ for GWCs. By combining the Hall conductivity $\sigma_{xy}$ and the charge distribution results, we construct the phase diagram in Fig. 2a: For $10\leq\epsilon_{r}\leq 15.5$, the system resides in a topologically trivial GWC phase with $\sigma_{xy}=0$; in contrast, within the range $15.5<\epsilon_{r}\leq 23$, the ground state transitions to a uniform FCI state with a quantized Hall conductivity of $\sigma_{xy}=1/3$.
III.3 Hole crystal for for $\nu=2/3$
@@
Figure 3: Emergence of QAHC states for $\nu=2/3$ filling. a Phase diagram containing CO, FCI, and QAHC phases versus $\epsilon_{r}$, characterized by $\rho(\bm{q})$ and $\sigma_{xy}$. b Charge density $n(\bm{r})$ and c hole density $h(\bm{r})$ distributions in the QAHC state obtained with $\epsilon_{r}=18$ and on $N_{y}=6$ cylinder. As we define an average density of $n=0.5$ (per site) as the “full” filling $\nu=1$ (per unit cell) of the CI phase, the hole filling is thus $h(\bm{r})=0.5-n(\bm{r})$, as depicted by green dots. The pink lines in panel c show the absolute values of $K^{ij}_{1}$ (see definition in the main text).
@@
In Fig. 3, we show the results for fractional filling $\nu=2/3$. Our scanning reveals that the ground state exhibits FMz order for $\leq\epsilon_{r}$ betwee 10 and at least 30, for $N_{y}=4,5,6$ cylinders. In Fig. 3a, we present the results of $\rho(\bm{q})$ and charge pumping for various $\epsilon_{r}$ and widths. Notably, for a system width of $N_{y}=6$, the system transitions from a CO state at relatively small $\epsilon_{r}<13$, to a uniform FCI state with maximum of $\rho(\bm{q})\simeq 0$ and $\sigma_{xy}=2/3$ for $13\leq\epsilon_{r}\leq 16.5$. Furthermore, for $\epsilon_{r}>16.5$ (and up to about $\epsilon_{r}\approx 30$), a QAHC state emerges with Hall conductivity of $\sigma_{xy}=1$ and an underlying GWC order of holes (see Fig. 3).
@@
We emphasize that the QAHC state exhibits strong geometric sensitivity. For the $N_{y}=6$ cylinder, we observed a quantized Hall conductivity of $\sigma_{xy}=1$; however, such a state is absent for $N_{y}=4$ and $N_{y}=5$. This can be understood by examining the underlying charge distributions. In Fig. 3b, we show real-space charge distribution $n(\bm{r})$, where the maximum density $n\simeq 0.470$ and minimum density close to zero. In Fig. 3c, the holes form a triangular GWC structure with a period of 3 along the width ($\mathbf{a}_{2}$) direction, which is compatible with the width $N_{y}=6$.
@@
Figure 4: FCI and QAHC for $\nu=1/2$ filling. a The quantized $\sigma_{xy}$ results obtained from charge pumping for various $\epsilon_{r}$ and $N_{y}$. b Charge pumping $|\Delta Q|$ after inserting flux $\Phi_{y}=4\pi$, for FCI and QAHC states with different $\epsilon_{r}$ and $N_{y}$. Panels c,e show $n(\bm{r})$ and $\rho(\bm{q})$, respectively, revealing non-uniform charge distribution for the FCI state obtained with $N_{y}=5$ and $\epsilon_{r}=12$. Panels d, f show results of the QAHC state for $N_{y}=6$ and $\epsilon_{r}=20$.
III.4 QAHC with band folding for for $\nu=1/2$
Besides the fractional fillings $\nu=p/q$ with odd denominators $q$ discussed above, we also consider an even-$q$ filling fraction $\nu=1/2$ in this study. As shown in Fig. 4b, charge pumping simulations on systems with an odd width $N_{y}=5$ (with $\epsilon_{r}=12$) find a fractional Hall conductivity $\sigma_{xy}=1/2$, indicating the emergence of a FCI state. However, the same simulations on even-width systems with $N_{y}=4,6$ (with a different $\epsilon_{r}=18$, though) find an integer Hall conductivity $\sigma_{xy}=1$, indicating the presence of QAHC state.
Figure 5: Charge excitation spectra in CI, FCI and QAHC phases. The band-resolved spectral functions $A_{\rm uu}(\mathbb{k},\omega)$ and $A_{\rm ll}(\mathbb{k},\omega)$ simulated at zero temperature are shown for a CI state at $\nu=1$, b FCI state at $\nu=2/3$, and c the QAHC state at $\nu=1/2$. The followed path in the reciprocal space is indicated in the inset of a. As a comparison, the bare band dispersion (without interactions) is illustrated by the dashed lines in a. The gray shadow in c represents the band folding with vector $\bm{q}_{\rm CO}=(0,-\pi)$. The peaks of the two folded bands are marked by dashed lines (serving as a guide for the eye). The frequency resolution is approximately 2.67 meV, as indicated by the range bar in a.
To understand such a width dependence, we compare the COs in the ground states. In Fig. 4, we show that for $N_{y}=5$, the charge distribution $n(\bm{r})$ is incommensurate with the honeycomb lattice and $\rho(\bm{q})$ shows peaks at $(0.866\pi,0.4\pi)$ and $(-0.866\pi,-0.4\pi)$. In contrast, for $N_{y}=6$, a well defined commensurate CO emerges, as evidenced by the $(0,\pm\pi)$ peaks of $\rho(\bm{q})$ in momentum space. For even $N_{y}$, the CO with an enlarged unit cell allows an integer filling of the folded band (see Fig. 5c and related discussions below), yielding integer Hall conductivity. Note that a recent ED study has also reported a QAHC state at $\nu=1/2$ filling but with a much larger symmetry-breaking pattern, at select $k$-points [24]. Here our large-scale simulation results further reveal a sensitive even-odd width dependence in the ground state. The emergence of a commensurate CO in even $N_{y}$ systems, and its absence for odd $N_{y}$, may explain the observations for $\nu=1/2$ filling. It offers an band-folding mechanism for the presence of QAHC, different from the hole crystal picture discussed above for $\nu=2/3$. However, we emphasize that these two scenarios — hole crystallization and band folding — are not mutually exclusive.

### v3

@@ III.1 Global phase diagram
The presence of FM order has been experimentally identified as a prerequisite for the emergence of FCI and other correlated topological states in MoTe2 [7, 10, 9, 8, 12, 11, 13]. Therefore, we first examine the emergence of FM order, characterized by full polarization of spin-$S_{z}$ (denoted as FMz). To evaluate the ground-state magnetism across various filling factors, we compare the energies of different $S_{z}$ sectors in our iDMRG simulations (see Supplementary section III online). Our results reveal a finite energy gap above the FMz ground state for $\nu\gtrsim 0.2$, as indicated by darker blue shading in Fig. 1d for $N_{y}=4$ and $5$. Additionally, for $N_{y}=6$, we find that FMz order remains robust at filling factors $\nu=1$, $2/3$, $1/2$, and $1/3$, consistent with recent experimental observations [7, 8].
@@
To investigate the presence of quantized Hall conductivity in the FMz phases, we employ a charge pumping method [69, 70, 71] within our iDMRG simulations. For filling factors $\nu=p/q$, we adiabatically introduce a flux of $\Phi_{y}=2q\pi$ along the cylinder axis by modulating the hopping terms with a phase factor $e^{i\Phi_{y}}c^{\dagger}_{i\sigma}c_{j\sigma}+h.c.$ as the electron traverses the periodic boundary. For a charge pumping of $\Delta Q$, the corresponding Hall conductivity is given by $\sigma_{xy}=\frac{2\pi|\Delta Q|}{\Phi_{y}}$, whose unit is set as $e^{2}/h$ in this study.
@@
Our simulations, primarily conducted for fillings $\nu=p/q$ and $q=1,2,3$, reveal the existence of quantized Hall conductivity $\sigma_{xy}$, as indicated in Fig. 1e. We also consider $q=5,7,9$ cases on $N_{y}=4,5$ cylinders, with the results summarized in Fig. S8 (Supplementary section III online). Notably, we successfully observe experimentally relevant $\sigma_{xy}=\nu=1$ for CI states (Supplementary section III online) and $\sigma_{xy}=\nu=2/3$, $4/7$, $5/9$, $4/9$, $3/5$ for FCIs. Intriguingly, we predict a robust FCI state with $\sigma_{xy}=\nu=1/3$ and advocate for its experimental observation in future studies. Moreover, we also identify instances where an integer Hall conductivity of $\sigma_{xy}=1$ can be observed at fractional fillings $\nu=1/2$, $2/3$ (for $N_{y}=6$ cylinder), as well as at $\nu=3/5$ and $4/5$ (for $N_{y}=5$). An even more exotic QAHC state emerges at an incommensurate filling $\nu\simeq 0.63$, as shown in Fig. S9 (Supplementary section III online). These integer Hall states coexist with COs that are compatible with the cylinder geometry, leading to the formation of QAHC states.
Figure 2: GWC and FCI phases for $\nu=1/3$ filling. (a) Phase diagram containing the GWC and FCI states tuned by the relative dielectric constant $\epsilon_{\mathrm{r}}$. The two phases are characterized by $\rho({\bm{q}})$ and the Hall conductivity $\sigma_{xy}$, obtained on $N_{y}=5$ and 66 cylinders.Charge distribution (b) $n(\bm{r})$ in real-space and (c) $\rho(\bm{q})$ in momentum space for the GWC state (with $\epsilon_{\mathrm{r}}=14$), showcasing the triangular lattice pattern of electron density modulation. (d) demonstrates $n(\bm{r})$ for the FCI state ($\epsilon_{\mathrm{r}}=18$), where a uniform charge distribution is observed. (e) Charge-charge correlation function $g^{nn}(\bm{r})$ in the FCI phase (with $\epsilon_{\mathrm{r}}=18$), displaying the charge distribution relative to the reference site $n_{0}$ marked by a black star. (f) Charge structure factor $S^{A}_{n}(\bm{q})$ in momentum space of the FCI state. (g) Charge pumping results $|\Delta Q|$ for various $\epsilon_{\mathrm{r}}$ for $N_{y}=6$, measured after inserting a flux $\Phi_{y}=6\pi$. The inset illustrates the charge pumping simulation process. For this and subsequent plots, the dashed hexagon [c.f., panels (c) and (f)] outlines the first Brillouin zones of the honeycomb lattice.
@@ III.2 GWC and FCI for $\nu=1/3$
To start, we consider the filling fraction $\nu=1/3$, and present the results on $N_{y}=5$ and $6$ cylinders. Our results reveal that the ground states exhibit FMz order for a broad range of $\epsilon_{\mathrm{r}}$ values, specifically $10\leq\epsilon_{\mathrm{r}}\leq 23$. For the regime with a relatively small $\epsilon_{\mathrm{r}}$, a GWC phase appear in the phase diagram Fig. 2a. The real-space charge density $n(\bm{r})$ is centered at the $A$ sublattice and forms a triangular lattice with primitive vectors $\bm{a}^{t}_{1}=(\sqrt{3},0)$ and $\bm{a}^{t}_{2}=(\sqrt{3}/2,3/2)$, thus breaking the translational symmetry of the original honeycomb lattice (see Fig. 2b). Correspondingly, in Fig. 2c we find clear peaks in $\rho(\bm{q})$ at the vertices ($\bm{K}$ points), namely, the characteristic reciprocal vectors $\bm{b}^{t}_{1}=(0,4\pi/3)$ and $\bm{b}^{t}_{2}=(-2\sqrt{3}\pi/3,2\pi/3)$ of the triangular lattice. The results confirm the presence of a triangular GWC for relatively small $\epsilon_{\mathrm{r}}$.
@@
However, for a larger relative dielectric constant, e.g., $\epsilon_{\mathrm{r}}=18$, the charge density $n(\bm{r})$ becomes uniformly distributed across the honeycomb lattice (Fig. 2d). As a result, $\rho(\bm{q})$ becomes negligible across the entire momentum space (Fig. 2a). As seen in Fig. 2a, the maximum of $\rho(\bm{q})$ undergoes a sharp drop to very small values at $\epsilon^{c}_{r}=15.5$. For even larger $\epsilon_{\mathrm{r}}>23$, the ground state no longer has an FMz order.
@@
To characterize the transitions between GWC and FCI, we compute the charge-charge correlations $g^{nn}(\bm{r})=\frac{\langle n_{0}n_{i}\rangle}{\langle n_{0}\rangle\langle n_{i}\rangle}$ between the charge density $n_{0}$ at a fixed site and other charge densities $n_{i}$ located at site $i$. This metric provides insights into the attractive ($g^{nn}>1$) and repulsive ($g^{nn}<1$) correlations between electrons. As depicted in Fig. 2e for $\epsilon_{\mathrm{r}}=18$, we find $g^{nn}<1$ for the first three nearest neighbors of the reference site, indicating a repulsive interaction among electrons within this close proximity. However, $g^{nn}$ exceeds 1.0 for the fourth and fifth nearest neighbors, suggesting an attractive correlation in this region. This oscillating pattern of repulsion and attraction can extend even further away from the reference site, which constitutes a characteristic of neutral collective excitation, roton, in a FCI liquid [72].
@@
which quantifies the spatial distribution of charge correlations in momentum space. In Fig. 2f, we plot $S^{A}_{n}(\bm{q})$ for $\epsilon_{\mathrm{r}}=18$ in the uniform state with $\rho^{m}(\bm{q})\simeq 0$. The $S^{A}_{q}(\bm{q})$ peaks at the $\bm{K}$ points, which reflect the oscillating pattern observed in $g^{nn}(\bm{r})$. The positions of the peaks align with those of the GWC peaks depicted in Fig. 2c, indicating that the condensation of roton modes may be responsible for the transition from a uniform FCI to a triangular-lattice GWC.
@@
Moreover, we simulate the Hall conductivity through charge pumping, as illustrated in Fig. 2g for several values of $\epsilon_{\mathrm{r}}$ and $N_{y}=6$, using iDMRG calculations. Our results demonstrate that inserting a flux of $\Phi_{y}=6\pi$ induces a charge pumping of $|\Delta Q|=1$ (thus $\sigma_{xy}=1/3$) for FCIs, while $|\Delta Q|=0$ for GWCs. By combining the Hall conductivity $\sigma_{xy}$ and the charge distribution results, we construct the phase diagram in Fig. 2a: For $10\leq\epsilon_{\mathrm{r}}\leq 15.5$, the system resides in a topologically trivial GWC phase with $\sigma_{xy}=0$; in contrast, within the range $15.5<\epsilon_{\mathrm{r}}\leq 23$, the ground state transitions to a uniform FCI state with a quantized Hall conductivity of $\sigma_{xy}=1/3$.
III.3 Hole crystal for $\nu=2/3$
@@
Figure 3: Emergence of QAHC states for $\nu=2/3$ filling. (a) Phase diagram containing CO, FCI, and QAHC phases versus $\epsilon_{\mathrm{r}}$, characterized by $\rho(\bm{q})$ and $\sigma_{xy}$. (b) Charge density $n(\bm{r})$ and (c) hole density $h(\bm{r})$ distributions in the QAHC state obtained with $\epsilon_{\mathrm{r}}=18$ and on $N_{y}=6$ cylinder. As we define an average density of $n=0.5$ (per site) as the “full” filling $\nu=1$ (per unit cell) of the CI phase, the hole filling is thus $h(\bm{r})=0.5-n(\bm{r})$, as depicted by green dots. The pink lines in panel (c) show the absolute values of $K^{ij}_{1}$ (see definition in the main text).
@@
In Fig. 3, we show the results for fractional filling $\nu=2/3$. Our scanning reveals that the ground state exhibits FMz order for $\epsilon_{\mathrm{r}}$ betwee 10 and at least 30, for $N_{y}=4,5,6$ cylinders. In Fig. 3a, we present the results of $\rho(\bm{q})$ and charge pumping for various $\epsilon_{\mathrm{r}}$ and widths. Notably, for a system width of $N_{y}=6$, the system transitions from a CO state at relatively small $\epsilon_{\mathrm{r}}<13$, to a uniform FCI state with maximum of $\rho(\bm{q})\simeq 0$ and $\sigma_{xy}=2/3$ for $13\leq\epsilon_{\mathrm{r}}\leq 16.5$. Furthermore, for $\epsilon_{\mathrm{r}}>16.5$ (and up to about $\epsilon_{\mathrm{r}}\approx 30$), a QAHC state emerges with Hall conductivity of $\sigma_{xy}=1$ and an underlying GWC order formed by holes (see Fig. 3).
@@
We emphasize that the QAHC state exhibits strong geometric sensitivity. For the $N_{y}=6$ cylinder, we can observe a quantized Hall conductivity of $\sigma_{xy}=1$; however, such a state is absent for $N_{y}=4$ and $N_{y}=5$. This can be understood by examining the underlying charge distributions. In Fig. 3b, we show real-space charge distribution $n(\bm{r})$, where the maximum density $n\simeq 0.470$ and minimum density close to zero. In Fig. 3c, the holes form a triangular GWC structure with a period of 3 along the width ($\mathbf{a}_{2}$) direction, which is compatible with the width $N_{y}=6$.
Figure 4: FCI and QAHC for $\nu=1/2$ filling. (a) Charge pumping $|\Delta Q|$ after inserting flux $\Phi_{y}=4\pi$, for FCI and QAHC states with different $\epsilon_{\mathrm{r}}$ and $N_{y}$. (b) The quantized $\sigma_{xy}$ results obtained from charge pumping for various $\epsilon_{\mathrm{r}}$ and $N_{y}$. Panels (c,e) show $n(\bm{r})$ and $\rho(\bm{q})$, respectively, revealing non-uniform charge distribution for the FCI state obtained with $N_{y}=5$ and $\epsilon_{\mathrm{r}}=12$. Panels (d, f) show results of the QAHC state for $N_{y}=6$ and $\epsilon_{\mathrm{r}}=20$.
@@
Figure 5: Charge excitation spectra in CI, FCI and QAHC phases. The band-resolved spectral functions $A_{\rm uu}(\mathbb{k},\omega)$ and $A_{\rm ll}(\mathbb{k},\omega)$ simulated at zero temperature are shown for (a) CI state at $\nu=1$, (b) FCI state at $\nu=2/3$, and (c) the QAHC state at $\nu=1/2$. The followed path in the reciprocal space is indicated in the inset of (a). As a comparison, the bare band dispersion (without interactions) is illustrated by the dashed lines in (a). The gray shadow in (c) represents the band folding with vector $\bm{q}_{\rm CO}=(0,-\pi)$. The peaks of the two folded bands are marked by dashed lines (serving as a guide for the eye). The frequency resolution is approximately 2.67 meV, as indicated by the range bar in (a).
III.4 QAHC with band folding for $\nu=1/2$
Besides the fractional fillings $\nu=p/q$ with odd denominators $q$, below we also consider an even-$q$ filling fraction $\nu=1/2$ in this study. As shown in Fig. 4a, charge pumping simulations on systems with an odd width $N_{y}=5$ (with $\epsilon_{\mathrm{r}}=12$) find a fractional Hall conductivity $\sigma_{xy}=1/2$, indicating the emergence of a FCI state. However, the same simulations on even-width systems with $N_{y}=4,6$ (with a different $\epsilon_{\mathrm{r}}=18$, though) find an integer Hall conductivity $\sigma_{xy}=1$, indicating the presence of QAHC state.
To understand such a width dependence, we compare the COs in the ground states. In Fig. 4, we show that for $N_{y}=5$, the charge distribution $n(\bm{r})$ is incommensurate with the honeycomb lattice and $\rho(\bm{q})$ shows peaks at $(0.866\pi,0.4\pi)$ and $(-0.866\pi,-0.4\pi)$. In contrast, for $N_{y}=6$, a well defined commensurate CO emerges, as evidenced by the $(0,\pm\pi)$ peaks of $\rho(\bm{q})$ in momentum space. For even $N_{y}$, the CO with an enlarged unit cell allows an integer filling of the folded band (see Fig. 5 and related discussions below), yielding integer Hall conductivity. Note that a recent ED study has also reported a QAHC state at $\nu=1/2$ filling but with a much larger symmetry-breaking pattern in real space [25]. Here our large-scale simulation results further reveal a sensitive even-odd width dependence in the ground state. The emergence of a commensurate CO in even $N_{y}$ systems, and its absence for odd $N_{y}$, may explain the observations for $\nu=1/2$ filling. It offers an band-folding mechanism for the presence of QAHC, different from the hole crystal picture discussed above for $\nu=2/3$. However, we emphasize that these two scenarios — hole crystallization and band folding — are not mutually exclusive.